Given f (x) = –^x2 + x + 2 and g(x) = square root2x+3 , find the following:
a. f g(x), find the domain
b. f g(3)
c. f f(x), find the domain

Answer by stanbon(75887)

Given f (x) = –x^2 + x + 2 and g(x) = square root(2x+3) , find the following:
a. f g(x), find the domain
f[g(x)] = f[sqrt(2x+3)] = -(2x+3)+sqrt(2x+3)+2 = -2x+sqrt(2x+3)-1
Domain: All Real Numbers >= -3/2
-----------------------------------------
b. f g(3) = f[sqrt(9)] = f(3) = -9+3+2 = -4
------------------------------------------
c. f f(x), find the domain
f[f(x)] = f[-x^2+x+2] = -(-x^2+x+2)^2+(-x^2+x+2)+2
Domain: All Real Numbers
------------
Cheers,
Stan H.
==========